OpenStudy (anonymous):

Solving this system of equations x + 2y = 10. Is y = 5 - x/2 the end result before graphing?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

How would I graph it?

OpenStudy (anonymous):

I know the 5 is the y intercept but the - x/2 is throwing me off

hartnn (hartnn):

to get x-intercept, put y=0 so get x from 0=5-x/2

hartnn (hartnn):

or u can get x from x + 2y = 10, but putting y=0 here, easier

OpenStudy (anonymous):

Ok so x = 10, making our problem y = 5 - 5 right?

hartnn (hartnn):

x=10 is correct x-intercept, so now u have both the intercepts, can u graph the line ?

OpenStudy (anonymous):

Yes, how about 3x + 4y = 8... how exactly do I do this one?

hartnn (hartnn):

u know the procedure to get intercepts ? to get x-intercept, put y=0, find x to get y-intercept, put x=0, find y

hartnn (hartnn):

so whats the y-intercept of 3x + 4y = 8 put x=0 there

OpenStudy (anonymous):

y intercept is 2. x intercept is ... 8/3?

hartnn (hartnn):

both are correct

OpenStudy (anonymous):

Bleh I hate graphing fractions. Thanks for the help.

hartnn (hartnn):

hmm...welcome :)